OpenStudy (anonymous):

Solve for c. A=1/3[(a-b)+c] a. c = 3A-(a-b) b. c = 3A-a-b c. c = 3a+3b=A d. c = 3Aab Anyone who can tell me the answer and explain how they got it will get a medal. Thanks!

OpenStudy (anonymous):

Dude...come on...basic simplification....

OpenStudy (anonymous):

@tanmaykhandait You don't have to be rude. If you don't want to answer the question then just ignore it.

OpenStudy (anonymous):

You are looking to get the value for c. So we must get c to one side of the equation. But first lets get rid of the fraction (1/3) and make it a value of 1. How would we do that?

OpenStudy (anonymous):

Well I get that you would divide 1/3 by both sides and get 3A = (a-b) + c but after that I don't understand what to do.

OpenStudy (anonymous):

Great. now we have to get a-b to the left side of the equation. You don't need the parentheses so if we have 3A=a+b+c what do we do to get a and b to the left side?

OpenStudy (anonymous):

Sorry should be 3A = a-b+c...

OpenStudy (anonymous):

So it would be c = 3A-a+b?

OpenStudy (anonymous):

Wait it would be c = 3A - (a+b) right?

OpenStudy (anonymous):

Yes but that can also be c=3A-(a-b) do you see how that works as we distribute the minus sign over (a-b)...Good job

OpenStudy (anonymous):

Okay, thanks for your help!